OpenStudy (anonymous):

Solve the equation for y. 7x - 3y = 4

OpenStudy (anonymous):

Subtract 7x from both sides. What do you get?

OpenStudy (anonymous):

-3y=-3

OpenStudy (anonymous):

\[\huge \color{green}{\textbf{Welcome To Openstudy..}}\]

OpenStudy (anonymous):

Not quite, the left side of your equation is correct but. On the right side, if you subtract 7x, wouldn't you have 4 - 7x?

OpenStudy (anonymous):

can anyone help me

OpenStudy (anonymous):

Do like this: \[\large 7x - 3y - 7x = 4 - 7x\] Now solve this..

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

There's not point in us telling you the answer, but we can help you to understand how to do it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You only need the answer?? Sorry we can't give you the answer itself and you have to cooperate with us to get your answer... I have written the equation above.. What will you get by solving that??

OpenStudy (anonymous):

Do you understand that you would get -3y = 4 - 7x ?? after subtracting 7x from both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@trany11 did u want the value of y in terms of x ??

OpenStudy (anonymous):

so i add 4+-7x together

OpenStudy (anonymous):

Okay, so that's always the first step when you're solving for x or y, get it alone by subtracting the other term from both sides. From here, do you have a guess at what you should do next to get just y = something?

OpenStudy (anonymous):

You can 't add these.. See 4 and 7x are not like terms.. You can add 4 with 7, you can add 4x with 7x but you can't add or subtract 4 and 7x..

OpenStudy (anonymous):

-3y=-3x is dhat right

OpenStudy (anonymous):

Why are you solving right hand side?? Keep it as such you cannot solve this..

OpenStudy (anonymous):

Not quite, at this point when you have a number attached to the term you want to get alone, in this case "y". You want to divide the other side by the number attached. So in this case, you want to divide both sides by -3, what do you get?

OpenStudy (anonymous):

i got y=x-7x

OpenStudy (anonymous):

again, your left side is correct but the right side is not. You're dividing *both* sides by -3 not just the left side. And dividing 4 by negative -3 doesn't equal x, right? What do you have when you divide (4-7x) by -3.

OpenStudy (anonymous):

9x

OpenStudy (anonymous):

Also, if something has an x attached to it, think of that as an apple. If it has a y attached think of that as a banana. And if neither x or y is attached think of that as an orange.

OpenStudy (anonymous):

So you can't add an 2apple + 2orange and get 4apple. You just have 2apple and 2orange.

OpenStudy (anonymous):

likewise if you divide 2apple/2orange you just have 2apple/2orange. Basically, x+x = 2x but x + 2 = x+2 not 2x or 2x/2 = 2x/2 not 1

OpenStudy (anonymous):

so the answer is y=x+

OpenStudy (anonymous):

Anyway, you'd end up having \[y = (4-7x)/-3 = -[(4-7x)/3)]\]

OpenStudy (anonymous):

You should study combining like terms and such. Ask your teacher, you need to understand the basics before you can do later algebra.

OpenStudy (anonymous):

which one is that y=x+....y=x+...y+x-

OpenStudy (anonymous):

y = (4-7x)/-3

OpenStudy (anonymous):

/ = divided by

OpenStudy (anonymous):

heyy

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what is 23 x 45

OpenStudy (anonymous):

You should use this technique: http://www.math-mate.com/chapter2_5.shtml Seriously, you will get nowhere by just getting answers, you have to try to understand why things are the way they are.